KTeV Shift Summary for OWL 9/27/99 (Mon)

Yee Bob Hsiung (hsiung@fnal.gov)
Mon, 27 Sep 1999 07:02:03 -0500

Shift Summary for OWL shift, 09/27/99 (Mon)

On Shift: Jason Ladue, Yee Bob Hsiung

Hours of beam available at all intensities: 7.0 (out of 7.0)
Hours of beam available at nominal intensity: 7.0
Hours of E799 data (special runs not included): 6.8
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time & intensity:
Re-init pipeline, kfrends
========================================================================
Runs ended this shift:

14737 -- BEAM_799 --

Tapes brought to FCC this shift:

UP(A,B,C)031-033, UPS011 (Run 14736)

========================================================================
Accesses during shift: None.

========================================================================
Overview of shift:

E799 data taking at 9E12 with 10" Be common absorber.
(Trigger 10 & 11 had new trigger definition since run 14736,
see logbookXXI, pg298.)
Fixed invisible online histogram display for pink terminals.
Re-init pipeline twice to fix CsI bad capid channel.
Onsplit tape mount 3 times.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14737 00:00 Shift Begins. Run 14737 is in progress.

00:40 JasonL fixed the invisible colors for E799 Filter
Process by changing ktscreen.nml file for our
pink terminals (No Green Color).

55 00:57 ktevana error --> kfrend.

75 01:23 ValP yells mayday to wake up CsI expert:
CsI1-342 has -76% gain drift, bad capid 3 times
ped=0.
78 Re-init pipeline w/DBC to see it helps or not,
Wake up Ed (CsI oncall expert) to consult.
~81 01:36 CsI-1342 problem (ValP) finally goes away.

84 ktevana error - kfrend.

103 02:02 ValP yells again to wake up CsI expert:
CsI-1381 has -49% gain drift, bad capid 184 times
ped=0.
105 Re-init pipeline w/DBC fixed problem.

178 03:42 "time to panic..." CsI-686 stab 26.8%
goes away next spill.

04:55 Onsplit request UPF007 mounted on ktev1.

06:06 Onsplit request UPH003 mounted on ktev4.

06:30 Onsplit request UPM007 mounted on ktev2.
========================================================================
Follow up Notes on Previous Shift

See below.

========================================================================
Items to Follow up on Next Shift

Monday, DAY : PLEASE:

(1) put the old HP4Si toner cartridges (at least 4 in control room) into
the FNAL recycling system.

(2) get the "BAD" pieces of electronics which litter the area moved over
to PREP for service.